If cot A= 12/5 and A is acute find sin3A and cos3A
Pls solve this fr me

Grade:11

1 Answers

Ajay
7 years ago
since cot A = base/perpclr  we have base = 12 and perplr = 5
htpns =  sqrt(144 + 25) = 13
hence Sin A = perpn/hyptns = 5/13
Now Sin 3A = 3sin A – 4 sinA
  Sin 3A =  3*5/13 – 4 (5/13)= 2035/2197
 
Now you can find Cos 3A = sqrt(1-Sin2 3A)